# RALLY: Role-Adaptive LLM-Driven Yoked Navigation for Agentic UAV Swarms

### v1

@@ IV-B1 Overall Performance
Fig. 6 presents the average rewards [^1] Consistent with our previous works Rif. [2], due to the large negative values of interference penalty $R_{\text{e}}^{t}$, and collision penalty $R_{\text{c}}^{t}$, the reward defined in RALLY is generally negative as well. over $30$ test episodes and demonstrates the impact of consensus mechanisms on task completion. Particularly, RALLY attains the highest mean reward and the narrowest variance distribution, indicating minimal variance, fastest convergence, and consistent task completion across various episodes. In contrast, CIHRL behaves conservatively, yielding stable but modest rewards. CoNavGPT achieves higher average rewards than CIHRL, confirming the LLM’s strong environmental understanding and decision-making, but it lacks online exploration and can get stuck in local optima. DITTO slightly outperforms CIHRL by using LLM-based self-cognition for role and action selection, yet its greedy role choices and absence of reinforcement feedback lead to high variance and unstable consensus. These results demonstrate that RALLY’s integration of RL–based environmental feedback and LLM-driven semantic decision making effectively guides the multi-UAV swarm to reach robust consensus and execute high-quality collaboration in complex dynamic environments.
@@
Next, we study the convergence of the credit-based role assignment mechanism. Particularly, “RMIX” uses the MLP-based fusion network for joint utility estimation detailed in Section III-B, while “VDN” aggregates $Q_{i}$ via a simple weighted sum $Q_{\text{tot}}=\sum_{i=1}^{n}w_{i}Q_{i}(o_{i},k_{i})$. Fig. 8 presents the corresponding results. Although both methods converge, RMIX converges faster and yields more accurate cumulative return estimates. Fig. 8 shows the LoRA fine-tuning loss curve, while the validation starting from step $500$ indicates successful convergence.
IV-B2 Generalization
@@
As depicted in Fig. 9, we evaluate RALLY and CIHRL with changing swarm sizes from $8$, $9$, $10$ to $11$ on the DS-CEFC task. As the swarm grows beyond this training configuration, CIHRL’s performance degrades substantially. In other words, without retraining for larger formations, the involved agents fall into “habitual grouping” patterns that repeatedly form the learned clusters, preventing effective coverage of additional targets and leading to significant score losses. In contrast, RALLY, which encodes the maximum permitted formation size into its prompt, dynamically forges a split consensus to avoid excessive clustering. More importantly, RALLY preserves high scoring ability even as the swarm size increases to $9$, $10$, and $11$. This clear contrast underscores RALLY’s superior generalization to larger, unseen formations than CIHRL.
IV-C Performance Analysis of RALLY
IV-C1 Contribution of RMIX
@@
To illustrate how RMIX enhances LLM semantic decision making, we examine the differences between the initial intention and inferred consensus of Agent #6, as shown in Fig. 10. In the initial LLM-only phase, Agent #6 computes its distance to the adversary ($\approx 6.86$ m), and driven by “safe aggregation” and “maximal scoring”, greedily adopts the Commander role with target (-8,8). While this choice yields short-term points, it overlooks team coordination and can destabilize consensus under complex threats.   Therefore, the role-selected policy overrides this role to Coordinator, shifting the agent’s motive from individual dominance to supporting and scheduling. Receiving a neighbor’s intent (also $(-8,8)$), the LLM then recommends $(-8,8)$ again. Moreover, the LLM also issues explicit role alignment that maintains both proximity to the Executor and collaboration with the Commander. This refined decision reuses the initial safety assessment and integrates multi-party communication via role mapping, yielding a more holistic trade-off between individual scoring and team consistency. By correcting the LLM’s isolated “Commander” bias, RALLY preserves LLM’s semantic planning strengths while injecting MARL’s distributed division of responsibility and information fusion. Consequently, the two-stage output achieves superior policy stability and collaborative performance in the dynamic DS-CEFC task.
IV-C2 Impact of Role Number
@@
Fig. 11 illustrates RALLY’s reward distributions under four different role configurations, including single role (Executor), two roles (CommanderExecutor), three roles (CommanderCoordinatorExecutor), and four roles (CommanderCoordinatorExecutorDecoy), where the Decoy role is specifically designed to divert enemy attention. The single-role setup achieves the lowest mean reward ($\approx-3,000$), or severely limited performance due to the absence of task decomposition and limited exploration–coverage trade-off. Introducing a dual-role hierarchy (CommanderExecutor) yields a slight increase in mean reward but greatly enlarged whiskers, indicating that overreliance on the Commander’s decisions amplifies fluctuation and undermines group synergy. In contrast, the three-role configuration combines high average performance with smaller variance, demonstrating that introducing a Coordinator role effectively mediates the semantic planning benefits from the LLM while enhancing consistency and robustness through MARL’s exploration and credit-assignment. Unfortunately, adding a fourth Decoy role reduces average reward and inflates variance, suggesting that excessive role granularity raises coordination overhead and consensus costs, thereby detracting from overall effectiveness. Overall, the three-role (CommanderCoordinatorExecutor) architecture strikes the optimal balance between performance and stability within our two-phase LLM–MARL convergence framework, fully leveraging semantic decision making and reinforcement-driven exploration to achieve superior formation coverage and convergence stability.
@@
TABLE III: Running performance of mainstream LLMs on an NVIDIA RTX 4090.
@@
IV-C3 LLM Finetuning
Fig. 12 present the fine-tuning performance for RALLY and compare with other models. It can be observed from Fig. 12 that a non-fine-tuned Qwen2.5-$7$B base model delivers markedly lower performance, while due to its occasional illegal outputs (e.g., “I suggest going to target point #8, #8.” or “region 8”), a direct calling of GPT-4o for online interaction possibly result in significant latency and network instability issues, degrading the performance. By contrast, RALLY, which fine-tunes Qwen2.5-7B base model under the dataset $\mathcal{B}_{\mathrm{fil}}$, harmonizes the high-quality inference of API-GPT-4o with the efficiency and stability of a smaller model.
Next, we evaluate the performance after fine-tuning a LoRA-based family of Qwen2.5 models with varying parameter counts ($0.5$-B, $1.5$-B, $3$-B, and $7$-B) on the DS-CEFC task. Fig. 13 summarizes the post-fine-tuning performance and demonstrates substantial performance gains. Furthermore, Table III reports the Average Inference Time (AIT), Memory Footprint (MF), and Runtime Overhead (RO) after running these models on an NVIDIA RTX 4090. It can be found that a Qwen2.5-$1.5$-B version strikes the balance by delivering robust decision quality with minimal inference overhead.
@@ IV-C3 LLM Finetuning
Figure 15: Task overview in Gazebo Simulator for SITL.
@@
IV-D Software-In-The-Loop Validation
@@
V Conclusion and Future Works
This paper introduces RALLY, an advanced LLM-MARL-integrated framework for UAV swarm control that combines LLM-driven semantic reasoning with MARL-based exploration. By integrating autonomous intent understanding, dynamic role assignment, and decentralized consensus building, RALLY enables each UAV to interpret local observations, select functional roles, and collaboratively decide on navigation goals under communication constraints. We validate RALLY in the MPE simulation environment and a high-fidelity Gazebo-ROS-PX4 SITL platform, demonstrating superior task completion, collaborative effectiveness, and generalization compared to existing methods in the DS-CEFC scenario.
Looking forward, we plan to optimize the lightweight LLM for faster inference and reduced communication latency, and to explore multimodal fusion and theoretical guarantees for rapid semantic consensus in larger UAV swarms. These efforts will pave the way for deploying intelligent, collaborative UAV systems in complex, resource-constrained missions.

### v2

@@ 4.2.1 Overall Performance
Fig. 6 presents the average rewards [^1] Consistent with our previous works Ref. [2], due to the large negative values of interference penalty $R_{\text{e}}^{t}$, and collision penalty $R_{\text{c}}^{t}$, the reward defined in RALLY is generally negative as well. over $30$ test episodes and demonstrates the impact of consensus mechanisms on task completion. Particularly, RALLY attains the highest mean reward and the narrowest variance distribution, indicating minimal variance, fastest convergence, and consistent task completion across various episodes. In contrast, CIHRL behaves conservatively, yielding stable but modest rewards. CoNavGPT achieves higher average rewards than CIHRL, confirming the LLM’s strong environmental understanding and decision-making, but it lacks online exploration and can get stuck in local optima. DITTO slightly outperforms CIHRL by using LLM-based self-cognition for role and action selection, yet its greedy role choices and absence of reinforcement feedback lead to high variance and unstable consensus. These results demonstrate that RALLY’s integration of RL–based environmental feedback and LLM-driven semantic decision making effectively guides the multi-UAV swarm to reach robust consensus and execute high-quality collaboration in complex dynamic environments.
@@
Next, we study the convergence of the credit-based role assignment mechanism. Particularly, “RMIX” uses the MLP-based fusion network for joint utility estimation detailed in Section III.3.2, while “VDN” aggregates $Q_{i}$ via a simple weighted sum $Q_{\text{tot}}=\sum_{i=1}^{n}w_{i}Q_{i}(o_{i},k_{i})$. Fig. 8 presents the corresponding results. Although both methods converge, RMIX converges faster and yields more accurate cumulative return estimates. Fig. 8 shows the LoRA fine-tuning loss curve, while the validation starting from step $500$ indicates successful convergence.
4.2.2 Generalization
@@
As depicted in Fig. 9, we evaluate RALLY and CIHRL with changing swarm sizes from $8$, 99, $10$ to $11$ on the DS-CEFC task. As the swarm grows beyond this training configuration, CIHRL’s performance degrades substantially. In other words, without retraining for larger formations, the involved agents fall into “habitual grouping” patterns that repeatedly form the learned clusters, preventing effective coverage of additional targets and leading to significant score losses. In contrast, RALLY, which encodes the maximum permitted formation size into its prompt, dynamically forges a split consensus to avoid excessive clustering. More importantly, RALLY preserves high scoring ability even as the swarm size increases to $9$, $10$, and $11$. This clear contrast underscores RALLY’s superior generalization to larger, unseen formations than CIHRL.
We further evaluate RALLY across three different target area configurations: the original $3\times 3$ grid, a $2\times 4$ grid, and a $4\times 2$ grid. The target locations in each scenario are randomly generated, ensuring dynamic and diverse environments. As shown in Fig. 10, RALLY performs consistently well across all three scenarios, with no significant difference in reward performance, reinforcing its ability to adapt to varying environmental conditions.
4.3 Performance Analysis of RALLY
4.3.1 Contribution of RMIX
@@
To illustrate how RMIX enhances LLM semantic decision making, we examine the differences between the initial intention and inferred consensus of Agent #6, as shown in Fig. 11. In the initial LLM-only phase, Agent #6 computes its distance to the adversary ($\approx 6.86$ m), and driven by “safe aggregation” and “maximal scoring”, greedily adopts the Commander role with target (-8,8). While this choice yields short-term points, it overlooks team coordination and can destabilize consensus under complex threats.   Therefore, the role-selected policy overrides this role to Coordinator, shifting the agent’s motive from individual dominance to supporting and scheduling. Receiving a neighbor’s intent (also (8,8)(-8,8)), the LLM then recommends (8,8)(-8,8) again. Moreover, the LLM also issues explicit role alignment that maintains both proximity to the Executor and collaboration with the Commander. This refined decision reuses the initial safety assessment and integrates multi-party communication via role mapping, yielding a more holistic trade-off between individual scoring and team consistency. By correcting the LLM’s isolated “Commander” bias, RALLY preserves LLM’s semantic planning strengths while injecting MARL’s distributed division of responsibility and information fusion. Consequently, the two-stage output achieves superior policy stability and collaborative performance in the dynamic DS-CEFC task.
4.3.2 Impact of Role Number
Fig. 12 illustrates RALLY’s reward distributions under four different role configurations, including single role (Executor), two roles (CommanderExecutor), three roles (CommanderCoordinatorExecutor), and four roles (CommanderCoordinatorExecutorDecoy), where the Decoy role is specifically designed to divert enemy attention. The single-role setup achieves the lowest mean reward ($\approx-3,000$), or severely limited performance due to the absence of task decomposition and limited exploration–coverage trade-off. Introducing a dual-role hierarchy (CommanderExecutor) yields a slight increase in mean reward but greatly enlarged whiskers, indicating that overreliance on the Commander’s decisions amplifies fluctuation and undermines group synergy. In contrast, the three-role configuration combines high average performance with smaller variance, demonstrating that introducing a Coordinator role effectively mediates the semantic planning benefits from the LLM while enhancing consistency and robustness through MARL’s exploration and credit-assignment. Unfortunately, adding a fourth Decoy role reduces average reward and inflates variance, suggesting that excessive role granularity raises coordination overhead and consensus costs, thereby detracting from overall effectiveness.
@@ 4.3.2 Impact of Role Number
Overall, the three-role (CommanderCoordinatorExecutor) architecture strikes the optimal balance between performance and stability within our two-phase LLM–MARL convergence framework, fully leveraging semantic decision making and reinforcement-driven exploration to achieve superior formation coverage and convergence stability.
@@
4.3.3 LLM Finetuning
Fig. 13 present the fine-tuning performance for RALLY and compare with other models. It can be observed from Fig. 13 that a non-fine-tuned Qwen2.5-77B base model delivers markedly lower performance, while due to its occasional illegal outputs (e.g., “I suggest going to target point #8, #8.” or “region 8”), a direct calling of GPT-4o for online interaction possibly result in significant latency and network instability issues, degrading the performance. By contrast, RALLY, which fine-tunes Qwen2.5-7B base model under the dataset $\mathcal{B}_{\mathrm{fil}}$, harmonizes the high-quality inference of API-GPT-4o with the efficiency and stability of a smaller model.
Table 3: Running performance of mainstream LLMs on an NVIDIA RTX 4090.
@@ 4.3.3 LLM Finetuning
Next, we evaluate the performance after fine-tuning a LoRA-based family of Qwen2.5 models with varying parameter counts ($0.5$-B, $1.5$-B, 33-B, and $7$-B) on the DS-CEFC task. Fig. 14 summarizes the post-fine-tuning performance and demonstrates substantial performance gains. Furthermore, Table 3 reports the Average Inference Time (AIT), Memory Footprint (MF), and Runtime Overhead (RO) after running these models on an NVIDIA RTX 4090. It can be found that a Qwen2.5-$1.5$-B version strikes the balance by delivering robust decision quality with minimal inference overhead.
Figure 15: Task overview in Gazebo Simulator for SITL.
@@
4.4 Software-In-The-Loop Validation
@@
5 Conclusion and Future Works
This paper introduces RALLY, an advanced LLM-MARL-integrated framework for UAV swarm control that combines LLM-driven semantic reasoning with MARL-based exploration. By integrating autonomous intent understanding, dynamic role assignment, and decentralized consensus building, RALLY enables each UAV to interpret local observations, select functional roles, and collaboratively decide on navigation goals under communication constraints. We validate RALLY via the MPE simulation environment and a high-fidelity Gazebo-ROS-PX4 SITL platform, demonstrating its superior task completion, collaborative effectiveness, and generalization compared to existing methods in the DS-CEFC scenario.
Looking forward, we plan to optimize the lightweight LLM for faster inference and reduced communication latency, and enhance system robustness through more large-scale settings and advanced communication strategies. In addition, we will address the possible local optima issue in CoT reasoning by exploring test-time training strategies and diversifying reasoning paths to improve reasoning diversity, reduce convergence to suboptimal solutions, and enhance generalization. We also aim to investigate multimodal fusion and theoretical guarantees for rapid semantic consensus in larger UAV swarms. These efforts will pave the way for deploying intelligent, collaborative UAV systems in complex, resource-constrained missions.
Appendix
In the Appendix, we provide the detailed prompts in Fig. 18 and give the reasoning sensitivity in Fig. 19.
Figure 18: Part of the prompts used in RALLY.
Figure 19: Specific prompts and corresponding responses by different agents running their distributed models in parallel. For brevity, we have omitted the repeated structural prompts and CoT reasoning, while retaining only the key data-containing components of the input and output.